A Statistics professor wants to estimate the mean Math SAT score of incoming freshmen. It is known that Math SAT scores are normally distributed. In a random sample of 20 incoming freshmen, the mean Math SAT score was 527 with a standard deviation of 88. Construct a 90% confidence interval for the population mean. Use the appropriate formulas and show your work! 6.__________________________[8]
Answer #1

df = n-1 = 20 - 1 = 19

T score for 90% confidence interval = t0.05,19 = 1.729

Confidence interval =(\bar{X} \pm t_{0.05,19}*\frac{s}{\sqrt{n}})

                                =(527 \pm 1.729*\frac{88}{\sqrt{20}})

                                =(527 \pm 34.02)

                                = (492.98 , 561.02)